Due: Tuesday, January 29th 1. Please watch these videos on bonds, Bond Valuation (3:58), Bond Basics (7:04), and Calculating YTM in EXCEL and TI w Bionic Turtle (8:57), then answer the following questions: $1,000 par value bond, has 7.5 percent, and 10 years until maturity. Assuming semi-annual coupon payments: a. b. What is the bonds price? an annual coupon rate of 6 percent, an annual yield to maturity of What is the bonds coupon payment per period? c. If the bond were selling for $929, what would the annual yield-to-maturity be? d. If the bond were selling for $929, what would the effective yield-to-maturity if you reinvest coupon payments at 9 percent?
Answer #1

USing excel to calculate bond price, YTM and reinvestment YTM

a) Bond's Coupon Payment per Period= 60 (=6% *1000)
b) Bond's Price= $897.04 Using Excel Formula=PV(7.5%,10,-60,-1000)
c) YTM= 7.01% Using Excel Formula=RATE(10,60,-929,1000)
d) FV of Coupons $911.58
Par Value 1000
Total $1,911.58
YTM= 7.48% Using Excel Formula=RATE(10,0,-929,1911.58)

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well